OpenStudy (adajiamcneal):

write the functionrule g(x) after the given transformations of the graph of f(x)=3x reflection in the x-axis ; vertical compression by a factor of 1/6

OpenStudy (adajiamcneal):

@Ashleyisakitty

OpenStudy (anonymous):

well a reflection on the x-axis is going to be -f(x), and a vertical compression is just going to be (1/6)f(x)

OpenStudy (adajiamcneal):

soo besides making it f(x)=-3x what do i do?

OpenStudy (anonymous):

f(x)= (1/6)(-)(3x) = -1/2x

OpenStudy (adajiamcneal):

thank you :D

OpenStudy (anonymous):

YW :)

OpenStudy (anonymous):

just on a side note... when you are doing transformations be sure you pay attention to which ones you change the actual function f(x) and which ones you change only the x in the function